o Chapter 7 Probiem Set Has Disease No 920 3,830 4,750 Has Risk FactorYes Yes No Total 80 170 250 Total 1,000 4,000 5,000 In this sample of men aged 55-70: The odds of having the disease are The odds of having the disease if the man has the risk factor are The odds of having the disease if the man does not have the risk factor are The odds of having the disease are than when the man does not have the risk factor. . times greater when the man has the risk factor . In this sample of men aged 55-70 The risk of having the disease is The risk of having the disease if the man has the risk factor is . The risk of having the disease if the man does not have the risk factor is times more likely to have the disease than a man without A man with the risk factor is the risk factor
Answer #1

odds of having disease =250/4750=0.0526

odds of having disease if man has the risk factor =80/920 =0.0870

odds of having disease if man does not have the risk factor =170/3830 =0.0444

the odds of having the disease are 1.9691 times greater when,,,,,,,,,,,,,,,

risk of having the disease =250/5000 =0.05

risk of of having disease if man has the risk factor =80/1000 =0.08

risk of having disease if man does not have the risk factor =170/4000=0.0425

a man with risk factor is 1.8824 times more likely to have the,,,,,,,,